OpenStudy (anonymous):

using the unit circle what is equivalent to sec19°

OpenStudy (anonymous):

Um, think I would just use a calculator for that one...

Directrix (directrix):

sec(x) = csc(90-x) sec(19) = csc(90-19) sect (19) = csc(71)